A projectile is fired south at an angle of 30 degrees to the horizontal on latitude 60 degrees. Find the derivation to east or west. The initial velocity is 10m/s


1
Expert's answer
2021-02-10T07:15:12-0500

Explanations


  • Consider the figure attached


  • The trajectory of the projectile can be illustrated as shown. At the given latitude the horizontal level of the person firing the projectile is the tangent drawn to the circle at that point.
  • Since the velocity is very low, it behaves in a simple projectile motion on earth a shown.
  • Since it was initially directed to the south, it moves towards the south only.
  • Therefore, no deviation to east or west.
  • Further, maximum height = "\large \frac{u^2\sin^2\theta}{2g}=\frac{(10ms^{-1})^2\sin^230}{2\times9.8ms^{-2}}=\small 1.276m"

Range = "\large \frac{u\sin2\theta}{g}=\frac{10ms^{-1}\times\sin60}{9.8}=\small 0.442m"